OpenStudy (anonymous):

A charge with a value of 3.0 x 10^-5 C is located 3.0 cm from a charge with a value of 6.0 x 10^-5 C. Determine the distance from the larger charge to the point where the total electric field is zero To answer this question, would I use the Fe=kq1q2/r^2?

OpenStudy (anonymous):

no use \[E = \frac{ Kq }{ r ^{2} }\] and add them to get zero resultant .

OpenStudy (anonymous):

Ok I did and got 9.0x10^8

OpenStudy (anonymous):

 \[E _{t}= k(\frac{ 6 \times10^{-5}}{ x ^{2} }- \frac{ 3\times10^{-5} }{ (0.03-x)^{2} })\] Et = 0 make cross multiplication